if the length of a rectangle is 2 km longer then twice the width, then what is the length if the area is 60 sqare kilometer
Answer by Paul(988)

Let the width be x
Let the length be 2x+2
Area = 60km^2.
Equation:
(x)(2x+2)=60
2x^2+2x=60
2x^2+2x-60=0
x^2+x-30=0
Factor:
(x-5)(x+6)
x=5
2(5)+2=12
Hence, the width is 5km and the length is 12km.
Paul.